OpenStudy (anonymous):

how do you find the equivalent fraction of 1 3/5 the top part of ----- over /25

OpenStudy (anonymous):

3/5 = 1*3/5 = 5/5*3/5 = (5*3)/(5*5) = 15/25 so 1 3/5 is equal to 1 15/25 Can you see how I've done that?

OpenStudy (anonymous):

Thank you.